# Matrix harmonic analysis at high temperature via the Dirichlet process

### v1

@@ Abstract.
We investigate harmonic analysis of random matrices of large size with their Dyson indices going simultaneous to zero, that is in the high temperature limit. In this regime, the limiting empirical spectral distribution and the multivariate Bessel function/Heckman-Opdam hypergeometric function of the empirical spectral distribution are intimately related to the Markov-Krein correspondence. The uniqueness, existence and other properties of the Markov-Krein correspondence are studied using the theory of the Dirichlet process.
@@ 1.1. Overview
Harmonic analysis studies functions and measures through their spectral representations—most notably via Fourier transforms or characteristic functions. One of the key strengths of harmonic analysis is that all information about a measure (or distribution) is encoded in its characteristic function.
@@
A particular interest is the asymptotics of the rank-one spherical integral, when the size of the matrix $N\to\infty$. A celebrated result [4, 29, 28] establishes a profound link between its logarithmic asymptotics and $\mathcal{R}$-transform from free probability theory, that is, an analogue of the Fourier transform in the non-commutative setting. Explicitly, If the spectral measure (the normalised counting measure of the eigenvalues) of the matrix $A_{N}$ converges weakly to a deterministic measure $\rho$, then as $N\to\infty$, with certain analytic conditions one has,
$\frac{2}{N\beta}\log I_{N}^{(\beta)}\left(\frac{N\beta}{2}u;A_{N}\right)\to\int_{0}^{2u/\beta}\mathcal{R}_{\rho}(s)\mathrm{d}s,$ (1.2)
where $\mathcal{R}_{\rho}$ is the $\mathcal{R}$-transform of the measure $\rho$. For recent developments in this direction, please also see [8, 27, 34].
The previous result is shown only for $\beta=1,2$. While classically $\beta$ represents the symmetry of the matrix spaces (with $\beta=1,2$ corresponding to orthogonal/unitary ensembles), more recent approaches establish a continuous deformation framework to reflect the continuous nature of the inverse temperature. This has been introduced by using theory of Macdonald polynomials (see e.g. [42, 5, 26]). A specific reduction of the Macdonald polynomial is the multivariate Bessel function (see e.g. [18, 48, 16, 24, 14, 52]), acting as the continuous deformation for the spherical integral, which coincides with the spherical integral for $\beta=1,2$,
@@
Specifically, we would like to consider $\beta$ being an $N$-dependent parameter, i.e. $\beta=\beta_{N}$ (we will omit the subscript $N$ throughout the rest of the paper). It turns out that there are three regimes of $\beta$:
@@
where $\rho$ and $\rho^{(c)}$ satisfies the Markov-Krein correspondence (1.11). Such result has been partially studied in [45, 3], assuming the existence of all moments. In this work we will extend it to more general limiting measures including the ones with unbounded supports and possibly with heavy tails.
@@
$J_{N}^{(\beta)}(\vec{s};A_{N}):=\int_{\mathrm{U}_{\beta}(N)}\prod_{k=1}^{N}\det([UA_{N}U^{*}]_{k\times k})^{s_{j}-s_{j+1}}\,\mathrm{d}U,\quad\beta=1,2.$ (1.6)
@@
Apart from the additive case, we will also investigate the asymptotics of the rank-one Heckman-Opdam hypergeometric function, in the high-terperature limit. If the normalised counting measure of the vector $\vec{a}$ has a limiting measure $\rho$, its rank-one Heckman-Opdam hypergeometric function tends to the Mellin transform of the $\rho^{(c)}$, which appears to be the same measure as ion (1.4),
@@
The unforeseen appearance of the Markov-Krein correspondence in this problem leads us to investigate the theory of Dirichlet process, a well-studied random process intimately relating with the Dirichlet distribution. It originally arises from the studies of Bayesian non-parametric statistics [21, 9, 10, 11, 17, 41, 40]. The Markov-Krein correspondence itself appears extensively in studies in Markov moment problems [39, 36], in Rayleigh measures [36, 38, 6, 19, 20, 22, 23], in asymptotic freeness [2, 47], and its Fourier transform [15, 41, 53]. The relationship between Dirichlet process and Markov-Krein correspondence is also described in Section 2.
@@
  • to recognise the counter part of the asymptotics of the spherical integrals [28, 46, 33] in the high temperature case;
  • to emphasise the previously neglected importance of the Dirichlet process in studies of the Markov-Krein correspondence and the high temperature limits;
@@ 1.2. Preliminaries and main results
1.2.1. Probability measures and convergence
In this paper we consider Borel probability measures on the real line. Their supports can be either compact or non-compact. Let $\rho$ be a Borel probability measure on $\mathbb{R}$ and $c>0$. We assume a logarithmic tail control of the measure $\rho$, that is, $\rho\in\mathcal{V}$ with
$\mathcal{V}=\left\{\rho:\rho\text{ is a probability measure and }\int\log(1+x^{2})\rho(\mathrm{d}x)<+\infty\right\},$ (1.10)
@@ 1.2.1. Probability measures and convergence
$\int\frac{\rho^{(c)}(\mathrm{d}x)}{(z-x)^{c}}=\exp\left(-c\int\log(z-s)\rho(\mathrm{d}s)\right),\quad\text{for all }z\in\mathbb{C}\backslash\mathbb{R}$, (1.11)
@@
where $a_{1}^{(N)}<\ldots<a_{N}^{(N)}$ are deterministic points on the real line.
The sequence $\{\rho_{N}\}$ approximates a limiting measure $\rho$ in the large $N$ limit. We want to introduce a specific type of convergence. For $0<\eta<1$, we introduce the extended $\eta$-Wasserstein distance by
@@
Then we give the following assumption.
Assumption 1.1.
There exists $0<\eta<1$ such that the extended $\eta$-Wasserstein distance between $\rho_{N}$ and $\rho$ is small when $N\to\infty$.
@@
    The convergence of the extended $\eta$-Wasserstein distance between probability measures is a strictly stronger requirement than the weak convergence. The reason why we require such an assumption is, as implied in Lemma 3.4, such convergence implies a uniform convergence (away from the support) of the logarithmic integral (the $g$-function). The latter is the essential technical requirements needed for our derivations.
@@
    Since the extended $\eta$-Wasserstein distance is bounded by the Wasserstein distance, which is bounded above by the Prokhorov metric, we conclude that in the case all $\rho_{N}$ and $\rho$ are (uniformly) compactly supported, Assumption 1.1 is equivalent to a weak convergence.
@@ 1.2.3. Main results
The main results of this paper are the asymptotics of those rank-one symmetric functions in different regimes. For the multivariate Bessel function, we have the following.
@@
[…]
@@
    $\lim_{N\to\infty}\mathcal{B}_{\vec{a}^{(N)}}\left(u;N,\frac{\beta_{N}}{2}\right)=\int_{\mathbb{R}}e^{ux}\rho(\mathrm{d}x).$ (1.20)
@@
    High-temperature regime: for $\beta_{N}=\frac{2c}{N}+o(1/N)$, in either the case $c>1$, or in the case $\rho$ being compactly supported, we have
    $\lim_{N\to\infty}B_{\vec{a}^{(N)}}\left(u;N,\frac{\beta_{N}}{2}\right)=\int_{\mathbb{R}}e^{ux}\rho^{(c)}(\mathrm{d}x).$ (1.21)
@@
The following theorem is the counterpart for the Heckman-Opdam hypergeometric function.
@@
[…]
@@
    $\lim_{N\to\infty}\mathcal{F}_{\log\vec{a}^{(N)}}\left(u;N,\frac{\beta_{N}}{2}\right)=\int_{0}^{+\infty}x^{u}\rho(\mathrm{d}x).$ (1.23)
@@
    High-temperature regime: for $\beta_{N}=\frac{2c}{N}+o(1/N)$ and $\mathrm{Re}\,u\geq 1$, in either the case $c>1$, or in the case $\rho$ being compactly supported in $\mathbb{R}_{+}$, we have
    $\lim_{N\to\infty}\mathcal{F}_{\log\vec{a}^{(N)}}\left(u;N,\frac{\beta_{N}}{2}\right)=\int_{0}^{+\infty}x^{u}\rho^{(c)}(\mathrm{d}x),\quad\mathrm{Re}\,u\geq 1.$ (1.24)
@@
Remark 1.3 (The technical assumptions).
@@
    For both Theorem 1.1 and Theorem 1.2 in the high-temperature regime, the analytic assumptions that $c>1$ or $\rho$ being compactly supported are only technical. They assures the contour integral (2.69) (in the case $\mathrm{Re}\,u=0$) to be absolutely integrable, which is a key technical point to the proof. We believe the same result holds without those assumptions as implied partially by results in [3].
@@ 1.2.4. Strategy of the proofs
Step I. We will show that for probability measures $\rho^{(c)}$ and $\rho$ satisfying the Markov-Krein correspondence, there is an contour integral formula for the characteristic function of $\rho^{(c)}$, in terms of $\rho$ (Theorem 2.13). This is done by investigating the theory of Dirichlet process.
Step II. We show that on the other hand, the multivariate Bessel function has an integral expression (Theorem 3.1). This agrees with the expression above (Corollary 3.3)in the following sense: when $\rho_{N}$ is a discrete measure, for any $c=\beta N/2$, the integral formula between the measures $\rho_{N}^{(c)}$ and $\rho_{N}$ is exactly the same as the one for the multivariate Bessel function $\mathcal{B}_{\vec{a}}(s;N,\beta/2)$, where $\vec{a}=(a_{1},\ldots,a_{N})$ gives the point mass of the discrete measure $\rho_{N}$. This is also a finite $N$ and discrete version of (1.21). This theorem is shown by generalising the existing integral formula for the multivariate Bessel function in [14] using a contour deformation.
@@
$\mathcal{F}_{\log\vec{a}}\left(u;N,\frac{\beta_{N}}{2}\right)=\int_{0}^{+\infty}x^{u}\rho_{N}^{(c_{N})}(\mathrm{d}x)$ (1.26)

### v3

@@ Abstract.
We investigate harmonic analysis of random matrices of large size with their Dyson indices going simultaneous to zero, that is in the high temperature limit. In this regime, we show that the multivariate Bessel function/Heckman-Opdam hypergeometric function of the empirical spectral distribution converges to the Fourier/Mellin transform of a measure, which and the limiting empirical distribution are intimately related by the Markov-Krein correspondence. The uniqueness, existence and other properties of the Markov-Krein correspondence can be studied using the theory of the Dirichlet process.
@@ 1.1. Overview
Harmonic analysis studies functions and measures through their spectral representations, most notably via Fourier transforms or characteristic functions. One of the key strengths of harmonic analysis is that all information about a measure (or distribution) is encoded in its characteristic function.
@@
A particular interest is the asymptotics of the rank-one spherical integral, when the size of the matrix $N\to+\infty$. A celebrated result [8, 40, 39] establishes a profound link between its logarithmic asymptotics and $\mathcal{R}$-transform from free probability theory, that is, an analogue of the Fourier transform in the non-commutative setting. Explicitly, if the spectral measure (the normalised counting measure of the eigenvalues) of the matrix $A_{N}$ converges weakly to a deterministic measure $\rho$, then as $N\to\infty$, with certain analytic conditions one has,
$\frac{2}{N\beta}\log I_{N}^{(\beta)}\left(\frac{N\beta}{2}u;A_{N}\right)\to\int_{0}^{u}\mathcal{R}_{\rho}(s)\mathrm{d}s,$ (1.2)
where $\mathcal{R}_{\rho}$ is the $\mathcal{R}$-transform of the measure $\rho$. For recent developments in this direction see [6, 38, 45].
The previous result is shown only for $\beta=1,2$. While classically $\beta$ represents the symmetry of the matrix spaces (with $\beta=1,2$ corresponding to orthogonal/unitary ensembles), more recent viewpoints treat $\beta$ not merely as a discrete symmetry parameter but as a continuous deformation parameter, reflecting the continuous nature of the inverse temperature of an associated interacting particle system. Two complementary frameworks allow one to extend the theory from the classical Dyson indices to all $\beta>0$.
In the flavour of statistical mechanics, the first framework, called the $\beta$-ensemble, is to identify eigenvalues as a one-dimensional Coulomb gas model subject to a confining potential and pairwise logarithmic repulsion. In this interpretation the parameter $\beta$ plays the role of the strength of repulsion (or inverse temperature), interpolating continuously between the classical symmetry classes and enabling a thermodynamic viewpoint on spectral statistics. A canonical example is the Gaussian $\beta$ ensemble, which has a joint density
$p(\mathrm{d}\lambda_{1},\ldots,\mathrm{d}\lambda_{N})\propto\exp\left(\beta\left(\sum_{j<k}\log|\lambda_{j}-\lambda_{k}|-\sum_{j=1}^{N}\frac{x_{j}^{2}}{4}\right)\right).$ (1.3)
We refer the readers to [4, 30, 25, 11, 10] for an overview of the classical $\beta$-ensembles.
Within the same Coulomb–gas interpretation, it is curious to allow the interaction strength $\beta$ itself to vary with the system size. This leads to the study of $N$-dependent $\beta$-ensembles, in which $\beta=\beta_{N}$ is a prescribed sequence depending on $N$. Of particular interest is the regime $\beta=\beta_{N}=O(\frac{1}{N})$ known as the high temperature regime, where the repulsion becomes asymptotically weak relative to the confining potential. In this regime the repulsion is sufficiently weak that classical log-gas methods must be modified, and many features of the spectrum exhibit a crossover between independent-particle behaviour and strongly correlated Coulomb gas behaviour. Following the pioneering analysis of the high temperature Gaussian $\beta$-ensemble in [3], a series of works have investigated high temperature $\beta$-ensembles for more general potentials and related particle systems [1, 60, 22, 2, 32, 31, 55, 62, 63, 64].
Although the $\beta$-ensemble viewpoint motivates why continuous $\beta$ is natural, our main results rely on the second deformation framework—one does not deform the repulsion strength but instead deforms the spherical integral itself. Such framework has been introduced by using theory of Macdonald polynomials (see e.g. [54, 9, 37]). A specific reduction of the Macdonald polynomial is the multivariate Bessel function (see e.g. [26, 65, 23, 35, 19, 70]), acting as the continuous deformation for the spherical integral, which coincides with the spherical integral for $\beta=1,2$,
@@
Again, we would like to consider $\beta$ being an $N$-dependent parameter, i.e. $\beta=\beta_{N}$ (we will omit the subscript $N$ throughout the rest of the paper). It turns out that there are three regimes of $\beta$:
@@
where $\rho$ and $\rho^{(c)}$ satisfies the Markov-Krein correspondence (1.12). Such result has been partially studied in [60, 58, 7], assuming the existence of all moments. In this work we will extend it to more general limiting measures including the ones with unbounded supports and possibly with heavy tails. We also refer the readers to [58, 7, 64] for several compelling examples of random matrix ensembles in the high temperature limit.
@@
$J_{N}^{(\beta)}(\vec{s};A_{N}):=\int_{\mathrm{U}_{\beta}(N)}\prod_{k=1}^{N}\det([UA_{N}U^{*}]_{k\times k})^{s_{j}-s_{j+1}}\,\mathrm{d}U,\quad\beta=1,2,\quad s_{N+1}=0.$ (1.7)
@@
A parallel phenomenon occurs in the multiplicative setting, governed by the asymptotics of the rank-one Heckman-Opdam hypergeometric function in the high-temperature limit. If the normalised counting measure of the vector $\vec{a}$ has a limiting measure $\rho$, its rank-one Heckman-Opdam hypergeometric function tends to the Mellin transform of the $\rho^{(c)}$, which appears to be the same measure as ion (1.5),
@@
The unforeseen appearance of the Markov-Krein correspondence in this problem leads us to investigate the theory of Dirichlet process, a well-studied random process intimately relating with the Dirichlet distribution. It originally arises from the studies of Bayesian non-parametric statistics [29, 14, 15, 16, 24, 53, 52]. The Markov-Krein correspondence itself appears extensively in studies in Markov moment problems [51, 48], in Rayleigh measures [47, 48, 50, 12, 27, 28, 33, 34, 68], in asymptotic freeness [5, 61], and in properties of its Fourier transform [21, 53, 71]. The relationship between Dirichlet process and Markov-Krein correspondence is also described in Section 2.
@@
  • to recognise the counterpart of the asymptotics of the spherical integrals [39, 59, 44] in the high temperature regime;
  • to emphasise the importance of the Dirichlet process in studies of the Markov-Krein correspondence and the high temperature limits;
@@ 1.2. Preliminaries and main results
1.2.1. Probability measures and convergence assumptions
In this paper we consider Borel probability measures on the real line. Their supports can be non-compact. Let $\rho$ be a Borel probability measure on $\mathbb{R}$ and $c>0$. We assume a logarithmic tail control of the measure $\rho$, that is, $\rho\in\mathcal{V}$ with
$\mathcal{V}=\left\{\rho:\rho\text{ is a probability measure and }\int_{\mathbb{R}}\log(1+x^{2})\rho(\mathrm{d}x)<+\infty\right\},$ (1.11)
@@ 1.2.1. Probability measures and convergence assumptions
$\int\frac{\rho^{(c)}(\mathrm{d}x)}{(z-x)^{c}}=\exp\left(-c\int_{\mathbb{R}}\log(z-s)\rho(\mathrm{d}s)\right),\quad\text{for all }z\in\mathbb{C}\setminus\mathbb{R},$ (1.12)
@@
where $a_{1}^{(N)}<\ldots<a_{N}^{(N)}$ are points on the real line. They can be considered either deterministic or random; in the latter case the measure $\rho_{N}$ is a random measure.
The sequence $\{\rho_{N}\}$ approximates a limiting measure $\rho$ in the large $N$ limit. We want to introduce a specific type of convergence. For $0<\eta<1$, the extended $\eta$-Wasserstein distance is given by
@@
Then we give the following assumptions.
Assumption 1.1 (Convergence for a sequence of deterministic measures).
There exists $0<\eta<1$ such that the extended $\eta$-Wasserstein distance between $\rho_{N}$ and $\rho$ is small when $N\to+\infty$.
@@
    The convergence of the extended $\eta$-Wasserstein distance between probability measures is a strictly stronger requirement than the weak convergence. The reason why we require such an assumption is, as implied in Lemma 3.4, such convergence implies a uniform control (away from the support) of the logarithmic integral (the $g$-function). The latter is the essential technical requirements needed for our derivations.
@@
    Since the extended $\eta$-Wasserstein distance is bounded by the Wasserstein distance, which is bounded above by the Prokhorov metric, we conclude that in the case all $\rho_{N}$ and $\rho$ are uniformly compactly supported, Assumption 1.1 is equivalent to a weak convergence.
  3. (3)
    In [39] an assumption regarding a small Dudley distance between $\rho_{N}$ and $\rho$ has been considered. In the special case where $\rho_{N}$ and $\rho$ are uniformly compactly supported, the convergence in Dudley distance is equivalent to the weak convergence. We would like to point out the similarity between our assumption and the one in [39].
@@
Assumption 1.2 (Convergence for a sequence of random measures).
Let $\{\rho_{N}\}_{N}$ be a sequence of random measures, i.e. all $a_{j}^{(N)}$’s are now random points. There exists $0<\eta<1$ such that the extended $\eta$-Wasserstein distance between $\rho_{N}$ and $\rho$ satisfies
$\lim_{N\to+\infty}\mathbb{E}\,\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)=0,\qquad\sup_{N}\mathbb{E}\,\exp\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)<+\infty.$ (1.16)
Remark 1.2.
By the Markov inequality, the second requirement implies an exponential bound of the tail of $\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)$ as
$\mathbb{P}\left[\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)<y\right]\leq e^{-y}\sup_{N}\mathbb{E}\,\exp\mathrm{d}_{W_{\eta}}(\rho_{N},\rho)<+\infty$ (1.17)
for $y>0$.
@@ 1.2.2. Symmetric functions
We also refer the readers to Propositions 3.1 and 3.2 for integal representations for the two rank-one symmetric functions.
@@ 1.2.3. Main results
The main results of this paper are the asymptotics of those rank-one symmetric functions in different regimes. For the multivariate Bessel function, we have the following result.
@@
Let $\{\rho_{N}\}$ be a sequence of discrete probability measures given in (1.13), and $\rho$ be its limiting measure under the convergence specified by Assumption 1.1. We also assume that $\rho\in\mathcal{V}$, and that for some $u\in\mathbb{C}\setminus\{0\}$ one has
@@
    $\lim_{N\to+\infty}\mathcal{B}_{\vec{a}^{(N)}}\left(u;N,\frac{\beta_{N}}{2}\right)=\int_{\mathbb{R}}e^{ux}\rho(\mathrm{d}x).$ (1.23)
@@
    High-temperature regime: for $\beta_{N}=\frac{2c}{N}+o(1/N)$, in the case $c>1$, we have
    $\lim_{N\to+\infty}B_{\vec{a}^{(N)}}\left(u;N,\frac{\beta_{N}}{2}\right)=\int_{\mathbb{R}}e^{ux}\rho^{(c)}(\mathrm{d}x).$ (1.24)
@@
Theorem 1.2 (The Fourier transform; the random case).
Let $\{\rho_{N}\}$ be a sequence of random discrete probability measures given in (1.13), and $\rho$ be its deterministic limiting measure under the convergence specified by Assumption 1.2. Then in the high-temperature limit $\beta_{N}=\frac{2c}{N}+o(1/N)$ with $c>1$,
$\lim_{N\to+\infty}\mathbb{E}\,B_{\vec{a}^{(N)}}\left(it;N,\frac{\beta_{N}}{2}\right)=\int_{\mathbb{R}}e^{itx}\rho^{(c)}(\mathrm{d}x),\quad t\in\mathbb{R}\setminus\{0\},$ (1.25)
where $\rho^{(c)}$ and $\rho$ satisfies a Markov-Krein correspondence, and the expectation takes over the randomness of $\vec{a}^{(N)}$.
@@
The following theorem is the counterpart of Theorem 1.1 for the Heckman-Opdam hypergeometric function.
@@
Let $\{\rho_{N}\}$ be a sequence of discrete probability measures on $\mathbb{R}_{+}$ given in (1.13), and $\rho$ be its limiting measure on $\mathbb{R}_{+}$ under the convergence specified by Assumption 1.1. We also assume that $\rho\in\mathcal{V}$, and that for some $u\in\mathbb{C}\setminus\{0\}$
@@
    $\lim_{N\to+\infty}\mathcal{F}_{\log\vec{a}^{(N)}}\left(u;N,\frac{\beta_{N}}{2}\right)=\int_{0}^{+\infty}x^{u}\rho(\mathrm{d}x).$ (1.27)
@@
    High-temperature regime: for $\beta_{N}=\frac{2c}{N}+o(1/N)$ and $\mathrm{Re}\,u\geq 1$, in the case $c>1$ we have
    $\lim_{N\to+\infty}\mathcal{F}_{\log\vec{a}^{(N)}}\left(u;N,\frac{\beta_{N}}{2}\right)=\int_{0}^{+\infty}x^{u}\rho^{(c)}(\mathrm{d}x),\quad\mathrm{Re}\,u\geq 1.$ (1.28)
@@
Remark 1.4 (Technical conditions and generalisations).
@@
    For Theorem 1.1, Theorem 1.2 and Theorem 1.3 in the high-temperature regime, the analytic assumptions that $c>1$ is only technical. They assures the contour integral (2.29) (in the case $\mathrm{Re}\,u=0$) to be absolutely integrable, which is a key technical point to the proof. We believe the same result holds without those assumptions, as it is implied partially by results in [7].
@@
  3. (3)
    For Theorem 1.2 we take $u=it\in i\mathbb{R}$. We believe this setting can be generalised to $u\in\mathbb{C}\setminus\{0\}$ with a similar condition (1.22) given in Theorem 1.1. However, this will further requires that the supports of the measures $\rho_{N}$ are uniformly bounded at least from one side of the real axis. We decide to keep this case open for the sake of convenience.
  4. (4)
    We believe that Theorem 1.2 can be generalised to the classical regime. To do so it requires a more detailed analysis on the error term in Section 3.3, and relates the difference in Stieltjes transforms to the difference in Fourier transform.
  5. (5)
    We believe that Theorem 1.2 can also be generalised to the case of Mellin transform. However, our integral representation (2.38) requires the support of the measure to be bounded either away from infinity or away from zero (when one takes the inverse of the random variable), which in turn requires for case-by-case more detailed analysis of the difference between the Mellin transforms.
@@ 1.2.4. Strategy of the proofs
Step I. We will show that for probability measures $\rho^{(c)}$ and $\rho$ satisfying the Markov-Krein correspondence, there is a contour integral formula for the characteristic function of $\rho^{(c)}$, in terms of $\rho$ (Theorem 2.9). This is done by investigating the theory of Dirichlet process.
Step II. We show that on the other hand, the multivariate Bessel function has an integral expression (Proposition 3.1). This agrees with the expression above (Corollary 3.3)in the following sense: when $\rho_{N}$ is a discrete measure, for any $c=\beta N/2$, the integral formula between the measures $\rho_{N}^{(c)}$ and $\rho_{N}$ is exactly the same as the one for the multivariate Bessel function $\mathcal{B}_{\vec{a}}(s;N,\beta/2)$, where $\vec{a}=(a_{1},\ldots,a_{N})$ gives the point mass of the discrete measure $\rho_{N}$. This is also a finite $N$ and discrete version of (1.24). This theorem is shown by generalising the existing integral formula for the multivariate Bessel function in [19] using a contour deformation.
@@
$\mathcal{F}_{\log\vec{a}}\left(u;N,\frac{\beta_{N}}{2}\right)=\int_{0}^{+\infty}x^{u}\rho_{N}^{(c_{N})}(\mathrm{d}x).$ (1.30)
@@
Step VI. Finally, to show Theorem 1.2, one needs to investigate the difference between both sides, which has been obtained in Step IV. Applying Assumption 1.2 to the difference gives the result.